# HaluEval-Wild: Evaluating Hallucinations of Language Models in the Wild

### v1

@@ Abstract
Hallucinations pose a significant challenge to the reliability of large language models (LLMs) in critical domains. Recent benchmarks designed to assess LLM hallucinations within conventional NLP tasks, such as knowledge-intensive question answering (QA) and summarization, are insufficient for capturing the complexities of user-LLM interactions in dynamic, real-world settings. To address this gap, we introduce HaluEval-Wild, the first benchmark specifically designed to evaluate LLM hallucinations in the wild. We meticulously collect challenging (adversarially filtered by Alpaca) user queries from existing real-world user-LLM interaction datasets, including ShareGPT, to evaluate the hallucination rates of various LLMs. Upon analyzing the collected queries, we categorize them into five distinct types, which enables a fine-grained analysis of the types of hallucinations LLMs exhibit, and synthesize the reference answers with the powerful GPT-4 model and retrieval-augmented generation (RAG). Our benchmark offers a novel approach towards enhancing our comprehension and improvement of LLM reliability in scenarios reflective of real-world interactions.
@@
Zhiying Zhu Carnegie Mellon University zhiyingz@cs.cmu.edu                        Zhiqing Sun Carnegie Mellon University zhiqings@cs.cmu.edu                        Yiming Yang Carnegie Mellon University yiming@cs.cmu.edu
@@ 2 Related Works
Recent studies have highlighted that language models often possess an awareness of their own knowledge (Kadavath et al., 2022), and the internal states of LLMs can recognize when they are producing misinformation (Azaria and Mitchell, 2023). These insights suggest that utilizing LLMs’ internal knowledge may offer a pathway to mitigate hallucinations. Several strategies have been proposed to enhance the factuality of LLM outputs. Sun et al. (2022) introduced a recitation mechanism, while Li et al. (2023b), Zou et al. (2023), and Chen et al. (2023b) focused on inference-time interventions.
@@
Benchmark HaluEval-Wild Avg. normal-$\downarrow$ MT-bench (score) normal-$\uparrow$ AlpacaEval normal-$\uparrow$ AlpacaEval 2.0 normal-$\uparrow$
@@
[…]
Query Type Direct Self-Reflection Hinted Self-Reflection Subcategory OoS CR IC BM CE All OoS CR IC BM CE All OoS CR IC BM CE All Llama 2-Chat 70B 64.00 83.00 34.69 70.71 49.00 60.28 54.00 78.00 28.57 67.68 41.00 53.85 39.00 74.00 16.33 35.36 22.00 37.34 Llama 2-Chat 13B 48.00 71.72 57.73 61.62 35.00 54.81 45.00 61.62 41.24 52.53 31.00 46.28 37.00 58.16 34.74 48.48 26.00 40.88 Llama 2-Chat 7B 54.00 73.00 57.73 64.65 33.00 56.48 49.00 63.00 30.93 56.57 31.00 46.10 45.00 56.00 16.48 56.57 24.00 39.61
Table 3: Hallucination rates of direct generation, self-reflection, and hinted self-reflection.
@@ 3 Construction of HaluEval-Wild
Real-user queries are vital for assessing LLM hallucination in practical scenarios. In this context, we introduce HaluEval-Wild, a challenging dataset curated from real-world interactions between individuals and LLMs. The construction pipeline of HaluEval-Wild is shown in Appendix A.
@@ 3.2 Fine-grained Categorization
We analyze query-induced hallucinations using a categorization framework, outlining five main types with examples in Appendix B:
@@
In our investigation, we instruct GPT-4 to automatically categorize 6,505 challenging queries labeled in the previous step (Section 3.1) into the aforementioned five fine-grained categories. The distribution of the five categories is detailed in Appendix E. However, the precision of the GPT-4 classifier may be compromised due to its inherent inclination towards hallucination. To enhance precision, we undertake manual verification for queries categorized under each type, retaining only those accurately classified and those in which Alpaca exhibits hallucinations. We conduct manual verification until each category reaches 100 instances. This meticulous validation not only confirms the potential for hallucination in such queries but also ensures that these queries pose sufficient challenges for LLMs, like Alpaca, to provide accurate answers. Ultimately, we establish a fine-grained benchmark for hallucination evaluation from user-LLM interactions in the wild.
@@ 3.3 Evaluation with Reference Answers
To facilitate the evaluation of hallucination in LLMs, we provide a reference answer generated by GPT-4 for each user query. To overcome the inherent hallucination challenges of GPT-4 and to provide a proficient response, we incorporate information from an external search engine [^2] https://duckduckgo.com/ by retrieving the top five relevant passages, which are then concatenated with the prompt for GPT-4. With the reference answer, we can evaluate an LLM response by asking GPT-4 to judge whether it is hallucinated. A response is considered non-hallucinatory if it is consistent with the reference answer or if GPT-4 explicitly admits its inability to fulfill the request. The prompts for automatic categorization, reference answer generation and hallucination evaluation are available in Appendix C.
@@ 4.3 Hallucination Mitigation with Self-Reflection
The hallucination rates of direct generation, self-reflection, and hinted self-reflection are illustrated in Table 3. There is a general trend of decreasing hallucination ratios when moving from direct generation to self-reflection, and further to hinted self-reflection, suggesting the effectiveness of self-reflection in reducing hallucination, especially with additional hints.
@@ 5 Conclusion
Appendix A Construction Pipeline
Figure 1: The construction pipeline of HaluEval-Wild.
Appendix B Fine-grained Challenging Query Examples with Hallucinated LLM Responses
Type Query Response
OoS Where am I located? You are located in the United States.
Table 4: Examples of various categories of challenging queries and their corresponding hallucinated LLM responses.
Appendix C Instructions for GPT-4
@@
Appendix D Benchmark Statistics
@@
Appendix E Query Types Distribution
In Figure  2, we illustrate the distribution of query types as determined by GPT-4. While acknowledging potential limitations in the precision of GPT-4’s classifications, the presented distribution provides valuable insights into the real-world prevalence of query types prone to inducing hallucinations in LLMs.

### v2

@@ Abstract
Hallucinations pose a significant challenge to the reliability of large language models (LLMs) in critical domains. Recent benchmarks designed to assess LLM hallucinations within conventional NLP tasks, such as knowledge-intensive question answering (QA) and summarization, are insufficient for capturing the complexities of user-LLM interactions in dynamic, real-world settings. To address this gap, we introduce HaluEval-Wild, the first benchmark specifically designed to evaluate LLM hallucinations in the wild. We meticulously collect challenging (adversarially filtered by Alpaca) user queries from existing real-world user-LLM interaction datasets, including ShareGPT, to evaluate the hallucination rates of various LLMs. Upon analyzing the collected queries, we categorize them into five distinct types, which enables a fine-grained analysis of the types of hallucinations LLMs exhibit, and synthesize the reference answers with the powerful GPT-4 model and retrieval-augmented generation (RAG). Our benchmark offers a novel approach towards enhancing our comprehension and improvement of LLM reliability in scenarios reflective of real-world interactions. Our benchmark is available at https://github.com/Dianezzy/HaluEval-Wild.
@@
Zhiying Zhu Carnegie Mellon University zhiyingz@cs.cmu.edu                        Yiming Yang Carnegie Mellon University yiming@cs.cmu.edu                        Zhiqing Sun Carnegie Mellon University zhiqings@cs.cmu.edu
@@ 2 Related Works
Recent studies have highlighted that language models often possess an awareness of their own knowledge (Kadavath et al., 2022), and the internal states of LLMs can recognize when they are producing misinformation (Azaria and Mitchell, 2023). These insights suggest that utilizing LLMs’ internal knowledge may offer a pathway to mitigate hallucinations. Several strategies have been proposed to enhance the factuality of LLM outputs. Sun et al. (2022) introduced a recitation mechanism, while Li et al. (2023b); Zou et al. (2023); Chen et al. (2023b) focused on inference-time interventions.
@@
Figure 1: The construction pipeline of HaluEval-Wild.
Table 1: Examples of various categories of challenging queries (OoS: Out-of-Scope Information, CR: Compelx Reasoning, IC: Inappropriate Content, BM: Beyond-Modality Interaction, CE: Confused / Errorneous Queries) and their corresponding hallucinated LLM responses.
Type Query Response
OoS Where am I located? You are located in the United States.
@@
Benchmark HaluEval-Wild Avg. $\downarrow$ MT-bench (score) $\uparrow$ AlpacaEval $\uparrow$ AlpacaEval 2.0 $\uparrow$
@@
Alpaca 7B 99.20%   4.54 26.46% 2.59%
Figure 3: Hallucination rates ($\downarrow$) of direct generation, SR (self-reflection), and hinted SR (hinted self-reflection.
@@ 3 Construction of HaluEval-Wild
Real-user queries are vital for assessing LLM hallucination in practical scenarios. In this context, we introduce HaluEval-Wild, a challenging dataset curated from real-world interactions between individuals and LLMs. The construction pipeline of HaluEval-Wild is shown in Appendix 1.
@@ 3.2 Fine-grained Categorization
We analyze query-induced hallucinations using a categorization framework, outlining five main types with examples in Appendix 1:
@@
In our investigation, we instruct GPT-4 to automatically categorize 6,505 challenging queries labeled in the previous step (Section 3.1) into the aforementioned five fine-grained categories. In Figure  2, we illustrate the distribution of query types as determined by GPT-4. While acknowledging potential limitations in the precision of GPT-4’s classifications, the presented distribution provides valuable insights into the real-world prevalence of query types prone to inducing hallucinations in LLMs.
To enhance the precision of the GPT-4 classifier, which may be compromised due to its inherent inclination towards hallucination, we undertake manual verification for queries categorized under each type, retaining only those accurately classified and those in which Alpaca exhibits hallucinations. We conduct manual verification until each category reaches 100 instances. This meticulous validation not only confirms the potential for hallucination in such queries but also ensures that these queries pose sufficient challenges for LLMs, like Alpaca, to provide accurate answers. Ultimately, we establish a fine-grained benchmark for hallucination evaluation from user-LLM interactions in the wild.
@@ 3.3 Evaluation with Reference Answers
To facilitate the evaluation of hallucination in LLMs, we provide a reference answer generated by GPT-4 for each user query. To overcome the inherent hallucination challenges of GPT-4 and to provide a proficient response, we incorporate information from an external search engine [^2] https://duckduckgo.com/ by retrieving the top five relevant passages, which are then concatenated with the prompt for GPT-4. With the reference answer, we can evaluate an LLM response by asking GPT-4 to judge whether it is hallucinated. A response is considered non-hallucinatory if it is consistent with the reference answer or if GPT-4 explicitly admits its inability to fulfill the request. The prompts for automatic categorization, reference answer generation and hallucination evaluation are available in Appendix A.
@@ 4.3 Hallucination Mitigation with Self-Reflection
The hallucination rates of direct generation, self-reflection, and hinted self-reflection are illustrated in Figure 3. There is a general trend of decreasing hallucination ratios when moving from direct generation to self-reflection, and further to hinted self-reflection, suggesting the effectiveness of self-reflection in reducing hallucination, especially with additional hints.
@@ 5 Conclusion
Appendix A Instructions for GPT-4
@@
Appendix B Benchmark Statistics